An object is at temp of 400 degree celcius.At what temp would it radiate energy twice as fast?The temp of the surroundings may be as sumed to be negligible

Grade:12

1 Answers

AskIITians Expert Hari Shankar IITD
14 years ago

By STefan-Boltzmann law, Rate of heat radiation E is proportional to T4 .

Let E = kT4.

Also, let the iniotial energy radiated be E1.

Temperature T = 400 C = 400+273 K = 673 K

So, E1 = k (673)4 .................(1)

Let rate of heat radiation be 2E1 at temperature T2.

Then, 2E1 = k(T2)4

2E1 = [E1/(673)4] (T2)4 using (1)

(T2)4 = 2 (673)4

T2 = (2)1/4 * 673 = 1.19 * 673 = 800.33 K